# Information-Theoretic Measures on Lattices for High-Order Interactions

### v1

@@ 6 Real-world applications
While traditional portfolio optimisation frameworks, which rely on limiting the covariance of investments [43], provide valuable insights into risk and return, they often overlook high-order interactions among stocks. Here, we compute the Streitberg information between stocks in the S&P 500 using their daily returns (assumed as $iid$ [44]) from 4 Jan 2010 to 24 Apr 2024. For comparison, we computed both the intra- and inter-sector pairwise, 3-, and 4-way Streitberg information.
As expected, results in Figure 4a shows that intra-sector pairwise interactions are significantly stronger than inter-sector (gray bar) across all sectors owing to, e.g., shared economic drivers or supply chain interdependencies. At 3-way we observe variation across sectors; sectors like Utilities which are heavily regulated exhibited increasing 3-way interactions, whilst sectors such as Health Care which are more diverse (including pharma companies, biotech firms and healthcare providers) displayed less 3-way interactions. We see further differences at 4-way Streitberg information. For example, the energy sector which is primarily driven by commodity prices, geopolitical events and regulations, had less 4-way interactions. On the other hand, Information Technology, a sector that is characterised by rapid innovation and interdependent product ecosystems, displayed more 4-way interactions (see Section F for more details).
We further computed the Streitberg information before and after 2 Jan 2020 to understand the impact of COVID on the interdependencies between stocks (Figure 4b). Pre-COVID, we found no significant difference in intra-sector 4-way interactions relative to inter-sector, however, since COVID we found that many sectors now show significantly more or less 4-way interactions relative to intra-sector.
Figure 5: High-order interactions improve decoding of trial stage and movement direction in neural activity during a centre-out reaching task. a Macaque performing repeated trials of a centre-out arm reaching task, with seven separate reaching directions. b Neural activity taken from the pre-frontal cortex during the preparation stage (knowledge of which direction to reach) and a go-stimulus after which the reaching action is made. c-d We computed the 2-, 3-, and 4-way Streitberg information between neurons at each time point, using trials as $iid$ samples. We then trained a logistic regression model to classify c reach-direction and d trial stage (preparation or action) using varying orders of interactions as input features. c Trial stage is best predicted using a combination of 2-, 3- and 4-way interactions, whilst d reach direction is readily decoded using just pairwise and 3-way high-order interactions. Paired t-test between accuracy of fitted models across sessions were significant (*, $p<0.05$) except where shown as (n.s.).
@@
Prior work has found higher-order interactions to play an important role in neuroscientific data [45, 46, 47, 48, 49, 15, 50]. Therefore, next, we asked whether high-order interactions improved decoding of electrophysiological recordings of a monkey performing a delayed centre-out hand-reaching task [51, 52]. In each recording session, the monkey was instructed to perform seven fixed hand movements for a randomly defined number of trials (Figure 5a). The spiking activity for each neuron is recorded using a 24-channel Utah array over a 1.2 second time period (500ms preparation period, 700ms action period, Figure 5b). The spiking data was converted to spiking rates using a Gaussian 20ms kernel. For a given session, we exhaustively computed the 2-, 3- and 4-way Streitberg information of all possible combinations across a set of 12 neurons. This was repeated for each time point of every condition across 12 sessions (more details in Section G).
First, we asked whether high-order interactions were more important during the preparatory stages or motor action. We trained a logistic classifier using varying orders of Streitberg information as features. To avoid overfitting due to the increasing number of features as the order increases, we applied PCA to the $d$-order interaction features and used a constant number of principal components. We found that individually 2-way Streitberg information was most informative, but by also including 3- and 4-way interactions the classification of the two stages was improved (Figure 5c), highlighting their role in neural coding [53]. Again, in the task of decoding reach direction, whilst individually the 2-way Streitberg information is most informative, highlighting the importance of redundancy in neural code [54], we found that a combination of 2- and 3-way Streitberg information is best for decoding reach direction (Figure 5d). The need for 4-way interactions to better predict preparation vs. action, but less useful for decoding reach direction, could be due to the need to integrate information for action planning [45]. In both cases, training a model only using $d$-order interactions led to reduced accuracy (Figure 12).
@@ 7 Discussion
In this work, we have established a framework that formulates high-order information-theoretic measures using lattice theory. Through the lens of our framework, we show that well-known information-theoretic measures are readily derived, and we introduce a novel information measure, Streitberg information, that effectively captures all factorisations of the joint distributions.
Our work offers a rigorous and systematic approach for the reconstruction of high-order systems using information-theoretic measures, but has some limitations. While the sign of $\mathrm{SI}(d)$ ($d\geq 2$) is determined by the difference between the total dependence and the sum of the lower order Streitberg information through the recursiveness property, how the sign changes with order $d$ remains unexplored.
Despite inherent differences between our measure and PID, whereby our measure is symmetric (not conditioned on a target set), we also show that our framework provides key insights into the formulation of PID via the free distributive lattice (Section H). We note that Streitberg information scales proportional to the Bell number, making it increasingly expensive for large $d$. However, this is considerably more efficient than PID which scales by the Dedekind number [55]. Moreover, the magnitude of Streitberg information is explained in terms of the factorisability of the joint probability distribution, however, the bounds of the measure across alpha remain unclear. Finally, the non-parametric estimation of the Streitberg information assumes that the data is $iid$, a strong assumption in many real-world data sets which contain time-series [39]. This leaves several open directions for future work, specifically the investigation of the sign and bounds of Streitberg information across order and alpha; how it can be accurately approximated when the data has temporal dependence; and if the lattice-based framework is helpful in formalising directed information-theoretic measures.
@@ Appendix A Proofs
A.1 Proof of Proposition 1
@@
where $\sum_{j}|S_{j}|=d$. This is known as the Pythagorean theorem of KL divergence in information geometry [36]. One can compute the coefficients of the resulting lower order KL divergence terms and will see cancellations occur with the original lower order terms. Instead here we show this can be easily achieved by the information measures defined on the interval lattices.
@@
The recursiveness of $\mathrm{SI}(d)$ and $\mathrm{LI}(d)$ in Lemma 2 shows that both measures can be expressed by a sum of themselves in lower order. If we take the difference between $\mathrm{SI}(d)$ and $\mathrm{LI}(d)$ we get:
@@
where $\mathrm{SI}(\pi_{s})$ is the Streitberg information defined on the interval lattice $[\hat{\mathbf{0}},\pi_{s}]$, a sublattice of $P(d)$ (for details of the measure on interval lattice, see Section B). Each $\mathrm{SI}(\pi_{s})$ is expressed by the sum of the KL divergences related to the factorisation of $\pi_{s}$ and equals zero due to the Pythagorean theorem above.
@@ A.3 Proof of Lemma 2
for all $\pi\in P(d)$. Moving the lower order sum to the left hand side we recover:
@@
Appendix B Generalised Streitberg information
@@
This provides a general framework to detect high-order information within blocks of variables. In fact, the generalised Streitberg information can be expressed in terms of ordinary Streitberg information in Definition 4 [27]:
@@
Appendix C Simplicial Complex
Appendix D Tsallis-Alpha Divergence for Multivariate Gaussian: Analytical Expression
@@ Appendix F Stock market data
Figure 11: $d$-order Streitberg information of stock returns before and after COVID. All the cleaned data before/after 2020 are used to compute the high-order information between stocks for the pre-COVID/COVID period.
For a given stock in the S&P 500, we calculate the daily return from 4 Jan 2010 to 24 Apr 2024 as the difference between its closing and opening prices on each day. The returns at different time points are generally assumed as $iid$ [44]. Next we drop the stocks that have more than 20% of missingness in time and then drop the time points if there is missingness in any stocks. As a result we have 2900 realisations of daily returns of 459 stocks in S&P 500. Each stock belongs to one of the 11 sectors as defined in Global Industry Classification Standard (GICS). We compute the pairwise, 3-, 4-way SI to the sets of stocks that are either taken from within the same sector or sampled at random from the 459 stocks. In each case, we sample 500 sets of regions or take all possible combinations, whichever is lowest.
@@ Appendix G Macaque reaching task data
Figure 12: Accuracy of macaque decoding models using $d$-order interactions separately. a Increasing order interactions contained less information for predicting preparation vs action stages. b Similarly, increasing order $d$ of interactions provided less predictive power for decoding reach direction
@@ Appendix I Computational considerations
The code and data used to perform the experiments in the paper are provided in the GitHub repository: https://github.com/neurips2024hoi/nips2024hoi.git.

### v3

@@ 6 Applications
While traditional portfolio optimisation frameworks, which rely on limiting the covariance of investments [47], provide valuable insights into risk and return, they often overlook higher-order interactions among stocks. Here, we compute the Streitberg information between stocks in the S&P 500 using daily returns (assumed as $iid$ [48]) from 4 Jan 2010 to 24 Apr 2024. For comparison, we computed both intra- and inter-sector pairwise, 3-, and 4-way Streitberg information (see Appendix F for details).
As expected, results in Figure 4a show that intra-sector pairwise interactions are significantly stronger than inter-sector (grey bar) across all sectors owing to, e.g., shared economic drivers or supply chain interdependencies. At 3-way, however, we observe variation across sectors; sectors like Utilities which are heavily regulated exhibited increased 3-way interactions, whilst sectors such as Health Care which are more diverse (including pharma companies, biotech firms and healthcare providers) displayed less 3-way interactions. We see further differences at 4-way Streitberg information. For example, the energy sector which is primarily driven by commodity prices, geopolitical events and regulations, had lower 4-way interactions. On the other hand, Information Technology, a sector that is characterised by rapid innovation and interdependent product ecosystems, displayed higher 4-way interactions. We further computed the Streitberg information before and after 2 Jan 2020 to study the impact of COVID on the interdependencies between stocks (Figure 4b). Pre-COVID, we found no significant difference in intra-sector 4-way interactions relative to inter-sector; however, since COVID almost all sectors now show significantly different 4-way interactions (mostly higher, some lower) relative to inter-sector.
Figure 5: Streitberg information significantly improves the decoding of preparatory stage vs. motor action in neural activity. Classifiers that include higher-order features from $\mathrm{SI}$ , $\mathrm{LI}$ and $\mathrm{TC}$ are trained to identify trial stage (preparation vs. action). We report the significance of paired t-tests between model accuracies across sessions (*, $p<0.05$; **, $p<0.005$; ns, not significant). Adding SI features ($d=3,4$) improves classifier accuracy, whereas adding higher-order LI and TC features induces no significant improvement in accuracy compared to the pairwise model (‘2’).
@@
Prior work has found higher-order interactions to play an important role in neuroscientific data [18, 49, 50]. Therefore, we asked whether higher-order interactions improved decoding of electrophysiological recordings of a monkey performing a delayed centre-out hand-reaching task [51, 52]. In each recording session, the monkey was instructed to perform seven fixed hand movements for a randomly defined number of trials. For a given session, we exhaustively computed the 2-, 3- and 4-way Streitberg information, Lancaster information and total correlation of all possible combinations across a set of 12 neurons. This was repeated for each time point of every condition across 12 sessions (details in Appendix G).
We asked whether higher-order interactions were more important during the preparatory stages or during motor action. We trained a logistic classifier using varying orders of $\mathrm{SI}$, $\mathrm{LI}$ and $\mathrm{TC}$ as features. To avoid overfitting due to the increasing number of features as the order increases, we applied PCA to the $d$-order interaction features and used a constant number of principal components for all orders. Our analysis revealed that while incorporating 3-way $\mathrm{TC}$ does not necessarily enhance accuracy, the inclusion of 3-way $\mathrm{SI}$, which is equal to $\mathrm{LI}$ for $d=3$, significantly boosts model accuracy. When we increased the order to 4, the Streitberg information once again improves the classification accuracy of trial stages, while the additional $\mathrm{TC}$ and $\mathrm{LI}$ fail, highlighting the importance of $\mathrm{SI}$ in neural coding.
Feature Selection
Feature selection is a critical step in machine learning, aiming to identify the most relevant variables that contribute to accurate predictions and model interpretability. Traditionally, feature selection techniques focus on assessing individual features based on their predictive power or statistical dependence with the target variable $Y$ [53]. However, such methods inherently overlook higher-order interactions between features—interactions that can be crucial in capturing the underlying structure of complex data sets—and can therefore lead to incorrect interpretations and feature relevance.
[…]
Methods $X_{1}$ $X_{2}$ $X_{3}$ $X_{4}$ $X_{5}$ MSE
[…]
In Table 1, we compare the performance of commonly used algorithms, such as recurrent feature selection [54] and SHAP [55], with methods based on higher-order information-theoretic measures. The set of GT features that explain $Y$ are exactly found with Streitberg Information, yielding the least mean squared error (MSE). Alternative methods select the wrong features since they are unable to identify the complex higher-order interactions between the input features and the target variable. $\mathrm{TC}$ and $\mathrm{LI}$ also fail as they cannot detect the partial factorisation in our GT. The methods that select less than three variables perform poorly, while the methods that select all five variables such as SHAP have comparably low MSE, but are both computationally more expensive and lack explainability since they select all features when only three should be selected.
@@ 7 Discussion
Our work offers a rigorous and systematic approach for the detection of higher-order interactions using information-theoretic measures, but has some limitations. While the sign of $\mathrm{SI}(d)$ is determined by the difference between the total dependence and the sum of the lower-order terms through the recursiveness property, how it changes with order $d$ remains unexplored. Moreover, although the magnitude of the Streitberg information can be related to the factorisability of the joint probability distribution, the bounds of the measure across $\alpha$ remain unclear.
A key motivation for asymmetric measures is their potential to provide insights into causality, but the link between our Streitberg information measure and causality would need further exploration. Despite the fundamental differences between our measure and PID, Streitberg information is more scalable compared with PID. This can be understood through our lattice-theoretical framework, which also provides theoretical insights into the formulation of PID via the free distributive lattice and its scaling with the number of variables (see Appendix H). This will be the focus of future work.
Other limitations include the fact that the non-parametric estimation of the Streitberg information assumes that the data is $iid$, a strong assumption in many real-world data sets, especially those that contain time series. Also the $k$NN-based estimation typically does not scale well with sample size or data dimensionality [56], and although recent work has explored using neural networks to estimate mutual information [57, 58], as of now, there are no similar estimators for Tsallis-Alpha divergence.
There are several open directions for future work. Specifically, the investigation of the sign and bounds of Streitberg information across order $d$ and $\alpha$; how $\mathrm{SI}$ can be accurately approximated when the data has temporal dependence; the connection between Streitberg information and causal discovery; linking the lattice-based framework with the formalisation of directed measures that split the variables into source and target sets; and defining more scalable and flexible estimators for Streitberg information.
Acknowledgements
MB acknowledges support by EPSRC through grant EP/W024020/1 funding the project “Statistical physics of cognition” and grant EP/N014529/1 funding the EPSRC Centre for Mathematics of Precision Healthcare at Imperial, and by the Nuffield Foundation under the project “The Future of Work and Well-being: The Pissarides Review”. RP acknowledges funding from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) Project-ID 424778381-TRR 295.
The authors would like to thank Pedro Mediano, Jianxiong Sun and Xing Liu for valuable discussions.
@@ Appendix A Proofs
A.1 Proof of Lemma 1
Proof.
A.2 Proof of Proposition 1
@@
where $\sum_{j}|S_{j}|=d$. This is known as the Pythagorean theorem of KL divergence in information geometry [38]. One can compute the coefficients of the resulting lower-order KL divergence terms and will see cancellations occur with the original lower-order terms. Instead here we show this can be easily achieved by the information measures defined on the interval lattices.
@@
The recursiveness of $\mathrm{SI}(d)$ and $\mathrm{LI}(d)$ in Lemma 2 shows that both measures can be expressed by a sum of themselves in lower-order. If we take the difference between $\mathrm{SI}(d)$ and $\mathrm{LI}(d)$ we get:
@@
where $\mathrm{SI}(\pi_{s})$ is the Streitberg information defined on the interval lattice $[\hat{\mathbf{0}},\pi_{s}]$, a sublattice of $P(d)$ (for details of the measure on interval lattice, see Section C). Each $\mathrm{SI}(\pi_{s})$ is expressed by the sum of the KL divergences related to the factorisation of $\pi_{s}$ and equals zero due to the Pythagorean theorem above.
@@ A.3 Proof of Lemma 2
for all $\pi\in P(d)$. Moving the lower-order sum to the left hand side we recover:
@@
Appendix B Simplicial Complex
Appendix C Generalised Streitberg information
@@
This provides a general framework to detect higher-order information within blocks of variables. In fact, the generalised Streitberg information can be expressed in terms of ordinary Streitberg information in Definition 4 [29]:
@@
Appendix D Additional synthetic experiments
Appendix E Tsallis-Alpha Divergence for Multivariate Gaussian: Analytical Expression
@@ Appendix F Stock market data
Figure 11: $d$-order Streitberg information of stock returns before and after COVID. All the cleaned data before/after 2020 are used to compute the higher-order information between stocks for the pre-COVID/COVID period.
For a given stock in the S&P 500, we calculate the daily return from 4 Jan 2010 to 24 Apr 2024 as the difference between its closing and opening prices on each day. The returns at different time points are generally assumed as $iid$ [48]. Next we drop the stocks that have more than 20% of missingness in time and then drop the time points if there is missingness in any stocks. As a result we have 2900 realisations of daily returns of 459 stocks in S&P 500. Each stock belongs to one of the 11 sectors as defined in Global Industry Classification Standard (GICS). We compute the pairwise, 3-, 4-way $\mathrm{SI}$ to the sets of stocks that are either taken from within the same sector or sampled at random from the 459 stocks. In each case, we sample 500 sets of regions or take all possible combinations, whichever is lowest.
@@ Appendix G Macaque reaching task data
Figure 12: Accuracy of macaque decoding models using $d$-order interactions separately. Increasing order interactions contained less information for predicting preparation vs action stages.
@@ Appendix I Computational considerations
The code and data used to perform the experiments in the paper are provided in the GitHub repository: https://github.com/barahona-research-group/streitberg-information.git.